OpenStudy (anonymous):

derivative of sin^2(x^2 +2)?

OpenStudy (anonymous):

use chain rule

OpenStudy (anonymous):

cos^2x(x^2+2)(2x) right?

OpenStudy (anonymous):

no 2sin(x^2+2)cos(x^2+2)(2x)

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

let ^2 be f let sin be g let x^2+2 be h, then d[f(g(h))]/dx = f'g'h'

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

so 3 inside each other? how would...? what.

OpenStudy (anonymous):

nvmind I will figure out tomorrow

OpenStudy (anonymous):

that's chain rule

OpenStudy (anonymous):

Let \[f(x)=x^2, g(x)=sinx, h(x)=x^2+2\] Then,\[\sin^2(x^2+2)=f(g(h(x)))\] Chain Rule, which is one of the ways to compute derivative is follows.\[a(b(x))'=a'(b(x))*b'(x)\] Applying chain rule twice for our problem makes it\[f(g(h(x)))'=f'(g(h(x)))*(g(h(x)))'=f'(g(h(x)))*[g'(h(x))*h'(x)]\] which is,\[4xsin(x^2+2)\cos(x^2+2)\]